• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name B. Quinn

    Log entry time 16:03:26 on November 8,2009

    Entry number 298360

    keyword=Day shift summary

    08:00 Start of shift, Both spectrometers are down with magnet problems.
    08:15 RC looked at magnet situation, decides we can't make use of beam.
    08:22 Told MCC we won't be using beam and confirmed we will make an access later. Meanwhile they will deliver ~5uA to give them enough beam in the machine for stability.
    09:02 RC paged Tech on call to deal with magnet problems. Requested hall be put in controlled access.
    09:40 Changed target to empty. Also hall survey is complete
    09:50 Tech, Gary, is here, went down to hall ~10:05
    10:20 Jack is here.
    11:53 Jack and Gary left hall. Closed hall. Right HRS is up, Left HRS is locked out.
    12:05 asked for harp scan. Posted values in elog are OK.
    12:18 Switch from Empty to holey Carbon with 6X6 raster.
    12:35 Going to LD2 with 2X2 raster, run 4863
    12:54 4X4 raster, run 4864
    13:02 6X6 raster, run 4864
    13:10 request 80-100uA Start production running
    13:18 Parity run 13814 (and right HRS run 4867)
    14:13 Parity run 13815 (and at 14:31, right HRS run 4868)
    14:50 Parity run 13816 (and at 15:32, right HRS run 4869)
    15:39 Parity run 13817
    15:58 Asked for escorted access to hall to look at DAQ problem on right arm
    16:02 End of shift ~3 hours of data taking with on HRS